# Robust Quantum Control via Multipath Interference for Thousandfold Phase Amplification in a Resonant Atom Interferometer

### v1

@@ Acknowledgements.
Acknowledgements. We thank Andre Carvalho, Garrett Louie, and Viktor Perunicic for valuable discussions, and Garrett Louie also for contributions to the apparatus. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. This work is funded in part by the Gordon and Betty Moore Foundation (Grant GBMF7945), the David and Lucile Packard Foundation (Fellowship for Science and Engineering), the Office of Naval Research (Grant Number N00014-19-1-2181), and the National Institute of Standards and Technology (Grant Number 60NANB19D168). This research was also supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. JG acknowledges support from a National Science Foundation (NSF) Quantum Information Science and Engineering Network (QISE-NET) Graduate Fellowship, funded by NSF award No. DMR-1747426. Author Contributions. The author list is grouped into authors who led one or more areas of work (names marked by a *), followed by additional authors, followed by the principal investigator at the end. Authors within each group are listed in reverse alphabetical order. TD led apparatus design and construction. KD, JG, KJ, SS, and YW contributed equally to data collection and analysis. KD and YW led development of the closed-loop optimization approach and spontaneous emission simulations and studies, the latter with supporting contributions from JG. JG (lead), ZC (supporting), KJ (supporting), and SS (supporting) developed the open-loop optimization approach. KD (equal), NS (equal), YW (equal), JG (supporting), KJ (supporting), and SS (supporting) commissioned the cold atom source and atom interferometer. TK (lead) and TD (supporting) supervised the work. All authors contributed to conceptualization of the work and validation of the results. KD (equal), JG (equal), KJ (equal), TK (equal), SS (equal), YW (equal), ZC (supporting), TD (supporting), and NS (supporting) prepared and reviewed the manuscript.

### v2

@@ Acknowledgements.
Acknowledgements. We thank Andre Carvalho, Garrett Louie, and Viktor Perunicic for valuable discussions, and Garrett Louie also for contributions to the apparatus. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. This work is funded in part by the Gordon and Betty Moore Foundation (Grant GBMF7945), the David and Lucile Packard Foundation (Fellowship for Science and Engineering), the Office of Naval Research (Grant Number N00014-19-1-2181), and the National Institute of Standards and Technology (Grant Number 60NANB19D168). This research was also supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology. JG acknowledges support from a National Science Foundation (NSF) Quantum Information Science and Engineering Network (QISE-NET) Graduate Fellowship, funded by NSF award No. DMR-1747426.
Author Contributions. The author list is grouped into authors who led one or more areas of work (names marked by a *), followed by additional authors, followed by the principal investigator at the end. Authors within each group are listed in reverse alphabetical order. TD led apparatus design and construction. KD, JG, KJ, SS, and YW contributed equally to data collection and analysis. KD and YW led development of the closed-loop optimization approach and spontaneous emission simulations and studies, the latter with supporting contributions from JG. JG (lead), ZC (supporting), KJ (supporting), and SS (supporting) developed the open-loop optimization approach. KD (equal), NS (equal), YW (equal), JG (supporting), KJ (supporting), and SS (supporting) commissioned the cold atom source and atom interferometer. TK (lead) and TD (supporting) supervised the work. All authors contributed to conceptualization of the work and validation of the results. KD (equal), JG (equal), KJ (equal), TK (equal), SS (equal), YW (equal), ZC (supporting), TD (supporting), and NS (supporting) prepared and reviewed the manuscript.
End Matter on Experimental Apparatus.—A horizontal laser beam [118] tuned to the 689 nm ${}^{1}S_{0}-^{3}P_{1}(m_{j}=0)$ intercombination line performs interferometry on a cloud of 88Sr atoms released from a 3D magneto-optical trap. The intermediate 21.6 $\mu$s lifetime of this transition is long enough to be suitable for clock atom interferometry and enables high (MHz-scale) Rabi frequencies with Watt-scale interferometer beams [74, 75], which is advantageous for measuring up to MHz-scale signals, such as in dark matter searches [4]. After an interferometer sequence, a 461 nm push beam spatially separates the ground state from the excited state; then the two populations are read out via fluorescence detection [74, 75]. The interferometer fringe visibility ($v$) and phase ($\Delta\phi$) are extracted by fitting the normalized excited state population ($P_{e}$) as a function of the final beamsplitter phase ($\phi_{B}$), to the expression $\frac{1}{2}(1+v\cos[\Delta\phi+\phi_{B}])$. An M-Labs ARTIQ timing system controls the power, phase, and duration of each pulse via an acousto-optic modulator. The ratio of the laser beam waist to the atom cloud size ($\approx 3$) and the ratio of the Doppler detuning spread to Rabi frequency ($\approx 0.1$, corresponding to a $\approx 4$ mK temperature) used in this work [113] are representative of the ratios anticipated for MAGIS-100.
For the $<200\;\mu$s sequence durations considered in this paper, the transverse displacement of the atoms in the beam during an experiment cycle due to free fall under gravity is small ($<200$ nm) compared to the waist of the interferometer beam (mm scale). For longer sequence durations, the interferometer beam can be oriented to be parallel to the direction of gravity to keep the interferometer centered on the atoms during an experiment cycle.
End Matter on Open-Loop Optimization Cost Function and Visualization.—The cost function can be tuned to adjust how strongly the population in stray trajectories is penalized as a function of distance from the central arms. Limiting the spread of trajectories could be useful for accommodating constraints on the interferometer size, like in cases where the interferometer needs to fit inside a satellite [36] or a magnetically shielded region [10]. The initial exploration of cost function optimization was done in part using the Q-CTRL Boulder Opal package [119]. Figure 4(b) presents a visualization of the spatial spread of the atoms for the UR-N sequences at large N, illustrating why these sequences do not perform well per our cost function. Figure 4(a) shows the corresponding reduction in phase sensitivity to resonantly oscillating signals for these sequences.
Figure 4: Comparison of UR-$N$ sequences [58] for different values of $N$. (a) Experimentally determined phase sensitivities for a 504-loop sequence. The normalized phase response is $(\Delta\phi[\delta\phi=\pi/(504\times 2)]-\Delta\phi[\delta\phi=0])/\pi$, where $\Delta\phi$ is the measured phase and $\delta\phi$ the amplitude of resonantly modulating phase. Error bars reflect uncertainties from sinusoidal fits. Panels (b), (c), and (d) show the spread in atom population for $N=16$, $N=32$, and $N=96$ in a 96-loop sequence, as determined by the semi-classical model.
End Matter on Resilience Against Detuning and Rabi Frequency Errors.—In Fig. 5, we examine the resilience of the optimized sequence to deviations in overall laser detuning and Rabi frequency by deliberately introducing errors in these parameters on the order of 10% of the nominal Rabi frequency. Despite these variations, the interferometer phase remains robust within approximately ±$\pm 100$ mrad and the visibility remains above 50% of its peak value.
Figure 5: Susceptibility of the $N=8$ optimized sequence to detuning (a) and Rabi frequency error (b) at 504 loops, as determined by experiment (solid points), and simulation (shaded bands, width indicating uncertainty). The interferometer visibilities (blue squares for experiment, upper shaded band for simulation) are normalized to their maximum value, and the interferometer phase is displayed in red circles with lower shaded band for simulation. Each data point is the average obtained from 5 individual fringe scans with the error bar being the standard deviation. The detuning is plotted relative to the estimated Rabi frequency for atoms at the center of the cloud [113].
End Matter on Closed-Loop Optimization.—In order to address experimental errors beyond effects captured by simulations, we survey the $\phi_{1}$-$\phi_{2}$ parameter space under the constraint described above by implementing closed-loop optimization, where optimization steps are determined by real-time experimental data [120, 121, 122, 123, 124]. We maximize a fitness function defined as $f=v\cos^{4}\left(\Delta\phi_{0}/2\right)$, where $v$ is the measured visibility and $\Delta\phi_{0}$ is the measured phase shift without an applied signal [113]. We begin the optimization near a point that our OBE simulations suggests maximizes the fitness function. The optimization trajectory converges to a pulse sequence consistent with simulation results and with the results of the open-loop optimization (Fig. 3(b)).